OpenStudy (anonymous):

if x=a cosθ - b sinθ and y=a sinθ + b cosθ, how would i show that x^2 +y^2=a^2+b^2

OpenStudy (alexwee123):

ugh trigonometry :/

OpenStudy (anonymous):

@alexwee123 right? tell me about it. i was doing fine until i met this strand in my course.

OpenStudy (asnaseer):

get an expression for x^2 get an expression for y^2 add them up collect common terms use the fact that \(\cos^2(\theta)+\sin^2(\theta)=1\) and answer to pop out...

OpenStudy (alexwee123):

ya i always hated it but i still got an A on my trig test :D

OpenStudy (anonymous):

@alexwee123 lucky.. ! i still have three more trig units after this -.-

OpenStudy (anonymous):

for x^2 i got acos^2θ + bsin^2θ , but for y^2 i got asin^2θ + absinθcosθ + absinθcosθ + bcos^2θ ....

OpenStudy (anonymous):

@asnaseer

OpenStudy (asnaseer):

your expressions are not correct, remember:\[(ab+cd)^2=a^2b^2+2abcd+c^2d^2\]and similarly:\[(ab-cd)^2=a^2b^2-2abcd+c^2d^2\]

OpenStudy (anonymous):

so (asinθ)^2 would be a^2sin^2θ right?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

k i'll try it again..

OpenStudy (asnaseer):

ok - good luck :)

OpenStudy (anonymous):

@asnaseer k so i did it and got \[x ^{2} = a^{2}\cos ^{2}(\theta)-2acos (\theta)bsin (\theta)+b ^{2}\sin ^{2}(\theta)\]\[y ^{2} = a^{2}\sin ^{2}(\theta)+2asin (\theta)bcos (\theta)+b ^{2}\cos ^{2}(\theta)\] then i know that \[[a^{2}\cos ^{2}(\theta)-2acos (\theta)bsin (\theta)+b ^{2}\sin ^{2}(\theta) ] + [a^{2}\cos ^{2}(\theta)-2acos (\theta)bsin (\theta)+b ^{2}\sin ^{2}(\theta)]\]\[= a^{2}\cos ^{2}(\theta)+b ^{2} \sin ^{2} (\theta)+a ^{2}\sin ^{2}(\theta) + b ^{2}\cos ^{2}\theta\] ... right?

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

you just need to collect "like" terms now and you are done

OpenStudy (asnaseer):

try and collect terms such that you end up with this expression somewhere:\[(cos^2(\theta)+\sin^2(\theta))\]

OpenStudy (asnaseer):

because that expression always equals one

OpenStudy (anonymous):

\[a^2\cos^2(θ)+b^2\cos^2(θ)+a^2\sin^2(θ)+b^2\sin^2(θ)\] ? :S

OpenStudy (asnaseer):

there are two terms here that have \(a^2\) in them and two terms that have \(b^2\) in them.

OpenStudy (asnaseer):

try and make use of that fact

OpenStudy (asnaseer):

e.g.\[a^2p^2+a^2q^2=a^2(p^2+q^2)\]

OpenStudy (anonymous):

\[a^2\cos^2(θ)+a^2\sin^2(θ)+b^2\sin^2(θ)+b^2\cos^2(θ)\]\[=a^2(\cos^2(θ)+\sin^2(θ))+b^2(\sin^2(θ)+\cos^2(θ))\]\[=a^2(1)+b^2(1)\]\[=a^2+b^2\] AIOSHADS ! :D YEAH ?

OpenStudy (asnaseer):

you got it! well done! :D

OpenStudy (anonymous):

yay ! thanks so much :)

OpenStudy (asnaseer):

yw - I'm glad I was able to help :)